OpenStudy (anonymous):

Determine if the Mean Value Theorem applies to the function on the given interval. If the either applies, find all values of c that are guaranteed by the theorem f(x)=x^3-x^2-2x on [-1,1] I do not understand, can some one show me the process?

ganeshie8 (ganeshie8):

wat does MVT say ?

ganeshie8 (ganeshie8):

First you need to show that the given function satisfies the hypothesis of MVT

OpenStudy (anonymous):

It states that if f(x) is defined and continuous on the interval [a,b] and differentiable on (a,b), then there is at least one number c in the interval (a,b) (that is a < c < b) such that \[f'(c)=\frac{ f(b)-f(a) }{ b-a }\]

ganeshie8 (ganeshie8):

Yes, since the given function is a simple polynomial, it is continuous in the given interval [-1, 1] and is also differentiable in (-1, 1). So it fits the hypothesis of MVT. we can apply MVT on this function

ganeshie8 (ganeshie8):

thats the first part : we have shown that f(x) is continuous + differentiable in given interval

OpenStudy (anonymous):

question: how do i do the process to find out if it's differentiable or continuous?

ganeshie8 (ganeshie8):

we dont have to prove it rigorously here, we just need show that it satisfies the conditions for MVT loosely

OpenStudy (anonymous):

oh ok

ganeshie8 (ganeshie8):

next, find f(b), f(a) and f'(c)

ganeshie8 (ganeshie8):

and plug them below and solve \(c\) \(\large f'(c)=\frac{ f(b)-f(a) }{ b-a }\)

OpenStudy (anonymous):

\[f'(c)=\frac{ f(1)-f(-1) }{ 1+1} \]

ganeshie8 (ganeshie8):

yes, evaluate f(1), f(-1) and f'(c) and plug them above

OpenStudy (anonymous):

will the answer be 1

ganeshie8 (ganeshie8):

\(\large f'(c)=\frac{ f(1)-f(-1) }{ 1+1}\) \(\large 3c^2 - 2c-2 =\frac{ -2 -0}{ 1+1}\) \(\large 3c^2 - 2c-2 =-1\) \(\large 3c^2 - 2c-1 = 0 \) \(\large (3c+1)(c-1) = 0\) \(\large c = -1/3 ~~~ c = 1\) discard 1 cuz \(1 \not \in (-1, 1) \) so oly one value for c : -1/3